# Towards AI Search Paradigm

### v1

@@ 6.1 Background
AI-powered search applications have experienced a substantial increase in user traffic. For example, prominent platforms such as Perplexity are reported to manage over 10 million daily visits (Perplexity, 2023). The core technology driving these applications consists of LLMs, which predominantly utilize the Transformer architecture (Vaswani et al., 2017). Although the Transformer model has demonstrated notable effectiveness, its computational complexity, which scales quadratically with the input sequence length, introduces significant challenges during inference. Moreover, in adherence to scaling laws (Kaplan et al., 2020), LLMs have evolved to encompass an exceptionally large number of parameters. For instance, models such as LLaMA-4, DeepSeek-V3, and Qwen-3 possess parameter counts ranging from 100 billion to 1,000 billion, thereby requiring considerable storage capacity.
The computational requirements of LLMs necessitate considerable hardware resources, and their inference latency constitutes a significant portion of the overall response time in AI search applications. For example, a state-of-the-art LLM for AI search, DeepSeek-R1 (DeepSeek-AI et al., 2025), which reportedly incurs high deployment costs (requiring 32 GPUs, each with 80 GiB memory per instance) and suffers from latency bottlenecks (a first token latency of 2.91 seconds and a decoding speed of approximately 30 tokens/second). Search applications are inherently latency-sensitive, as users typically have limited tolerance for waiting times.
@@ 6.2.1 Local Attention
The internal computations of LLMs are predominantly characterized by matrix multiplications involving input context tokens and model parameter matrices. A prominent strategy to mitigate this computational burden is to restrict the scope of input tokens attended to by each token, a technique termed local attention. It is well-established that the computational complexity of the standard full attention mechanism exhibits quadratic scaling ($O(N^{2})$) with respect to the input sequence length $N$. However, a substantial body of research (Xiao et al., 2023b) suggests that full attention is not always necessary, as only a subset of contextual tokens typically exerts a significant influence on the model’s output. Consequently, local attention mechanisms can often serve as an effective substitute for full attention.
@@
While various lightweight attention mechanisms have been proposed, such as linear attention (MiniMax et al., 2025), RWKV (Peng et al., 2023), and Mamba (Gu and Dao, 2024), their adoption typically requires training models de novo, incurring substantial costs. In contrast, local attention techniques can often be integrated into pre-trained full-attention models with minimal overhead. This adaptation can be accomplished either through lightweight fine-tuning (Zhang et al., 2025b) or, in some cases, applied directly at inference time without any retraining, as demonstrated by methods like StreamingLLM (Xiao et al., 2023b).
@@
Recent works have explored various strategies to adapt pre-trained models. For instance, LightTransfer (Zhang et al., 2025b) addresses the high computational cost of full attention by identifying and modifying “lazy layers” where attention patterns are highly localized. Its innovation lies in selectively replacing full attention in these layers with a more efficient mechanism like streaming attention. This approach significantly reduces the KV cache footprint and computational load while preserving long-context capabilities, and can be applied either training-free or with minimal fine-tuning.
@@
More advanced approaches aim to linearize pre-trained Transformers to achieve linear-time inference. A key challenge in this conversion is integrating recurrent mechanisms, such as gates, without costly retraining. Liger (Lan et al., 2025) tackles this problem by repurposing existing weights from the pre-trained model’s key projection to construct these gating mechanisms, thereby avoiding new parameters. Its hybrid attention, combining local sliding window attention with a linear recurrent model, effectively balances local context processing with long-range dependency modeling.
Similarly, LoLCATs (Zhang et al., 2025a) focuses on converting standard softmax attention into a subquadratic alternative, with a strong emphasis on scalability to extremely large models. The core innovation is a two-stage framework: an “Attention Transfer” step first trains a subquadratic mechanism to mimic the original attention, followed by a “Low-rank Linearizing” step that uses LoRA to correct residual errors. This parameter-efficient method has successfully linearized models as large as 405B parameters, achieving up to a 3$\times$ throughput increase while maintaining performance comparable to the original Transformer.
@@
Furthermore, models adapted with local attention can leverage existing, highly optimized attention libraries, such as FlashAttention (Dao et al., 2022), during deployment, thereby ensuring computational efficiency in practice.
@@ 6.2.2 Model Pruning
Recent works have developed various innovative approaches to structured pruning. For instance, Layer Collapse (LaCo) (Yang et al., 2024) introduces a training-free method that directly removes entire layers from a pre-trained model. Its core innovation is the Reserving-Differences-while-Seeking-Common (RDSC) Layer Merge principle, which calculates the parameter differences between adjacent layers and merges this differential into a subsequent layer. This “collapses” layers, effectively pruning them while preserving the model’s overall function. A key advantage of LaCo is its ability to remove 30% to 50% of layers while maintaining performance and, crucially, preserving the model’s internal architecture. This allows the pruned model to serve as a drop-in replacement in existing systems without code modifications.
Other methods focus on more granular structures. The CoFi (Coarse-to-Fine) pruning method (Xia et al., 2022) demonstrates that structured pruning can achieve compression and speedup comparable to distillation but at a much lower computational cost. CoFi’s innovation lies in simultaneously pruning both coarse-grained structures (e.g., layers) and fine-grained units (e.g., attention heads, hidden dimensions) using masks of varying granularity. To maintain accuracy, it integrates a dynamic layerwise distillation method that learns the optimal layer mapping between the teacher and the pruned student model, achieving over 95% sparsity and 10$\times$ speedups while retaining high accuracy.
Building on classic compression frameworks, SlimGPT (Ling et al., 2024) adapts the Optimal Brain Surgeon (OBS) technique for the unique challenges of structured pruning in LLMs. To overcome the limitations of OBS, SlimGPT introduces a Batched Greedy Pruning algorithm that enables low-cost and rapid structured pruning by processing groups of parameters. It employs a grouped Cholesky decomposition for attention heads and a dynamic group size for FFNs to achieve near-optimal pruning efficiently. A key advantage of SlimGPT is its accessibility: it is a task-agnostic method requiring only a small calibration dataset and can complete the entire pruning process in about an hour on a single GPU.
Another line of work targets semi-structured pruning, which is critical for hardware acceleration but can often degrade model accuracy. A novel plug-and-play post-training pruning (PTP) method (Zhang et al., 2024) addresses this by introducing two key components. First, a new metric, Relative Importance and Activation (RIA), jointly considers weight magnitude and activation data to prevent the inadvertent removal of important weight channels, a drawback of previous metrics. Second, it employs an efficient channel permutation algorithm that reorganizes input channels to maximally preserve important weights before enforcing the N:M sparsity pattern. This training-free approach is scalable to models over 70B parameters and makes sparse LLM inference more practical by converting models to a hardware-friendly N:M format with minimal performance loss.
@@ 6.3.1 Output Length Reduction
  • Prompt-based Length Reduction. Some techniques leverage prompting to curtail output length. For example, TALE-EP (Han et al., 2024) estimates token budgets by first prompting the LLM to determine a reasonable token limit for a given task, then incorporating this estimate into a constraint prompt that guides the model to generate concise yet accurate responses. Chain-of-Draft (CoD) (Xu et al., 2025) addresses the verbosity common in step-by-step reasoning by instructing models to “Think step by step, but only keep a minimum draft for each thinking step, with at most five words,” thereby preserving essential intermediate reasoning while significantly reducing token usage. Lee et al. (Lee et al., 2025) conducted a systematic study of different compression instructions, revealing a universal trade-off between reasoning length and accuracy, and proposing the concept of “token complexity” as the minimum number of tokens required for successful problem-solving.
@@
  • Training-based Length Reduction. Other methods achieve output length reduction through modifications during the training phase. Demystifying (Yeo et al., 2025) implements a Cosine Reward function based on Dirichlet principles, incorporating an “exceed length penalty” to control Chain-of-Thought (CoT) reasoning length while maintaining performance. Aggarwal et al. (Aggarwal and Welleck, 2025) modify the training data by incorporating explicit length constraint instructions, such as “Think for N tokens,” before applying policy optimization to pre-trained reasoning models. O1-Pruner (Luo et al., 2025) introduces a Length-Harmonizing Reward that calculates the ratio between reference model and predicted CoT lengths, incorporating accuracy constraints to prevent performance degradation while shortening reasoning processes. DAST (Shen et al., 2025a) employs SimPO to fine-tune models using a specially constructed length-preference dataset based on a self-defined token-length budget measurement. Kimi (Team et al., 2025) incorporates a length penalty into its policy optimization approach to improve long CoT activations and facilitate model merging.
@@
  • Compression of Intermediate States. A distinct category of techniques focuses on compressing intermediate data generated during the model’s inference process. For instance, Coconut (Chain of Continuous Thought) (Hao et al., 2024) treats final-layer hidden states as “continuous thought” to replace discrete tokens, reusing these hidden states as subsequent input embeddings and gradually adding latent CoT tokens during training, thereby improving accuracy while reducing intermediate “thinking” tokens. CODI (Shen et al., 2025b) employs self-distillation where the model serves as both teacher and student, jointly learning explicit and implicit CoT while aligning hidden activations, enabling internal reasoning without generating explicit tokens. CCOT (Cheng and Van Durme, 2024) condenses long reasoning into short, contentful contemplation tokens by precomputing full CoT, selecting important hidden states, and training LoRA modules to predict and decode these key tokens.
@@ 6.3.2 Semantic Caching
AI search applications frequently encounter a high degree of semantic similarity among queries submitted by different users. Consequently, a considerable fraction of distinct input queries can yield identical or substantially similar answers (e.g., “Tell me about the history of Hawaii” versus “What is the history of Hawaii?”). This redundancy is leveraged by various optimization strategies, including query caching, result caching, and query rewriting (Gill et al., 2025).
For instance, MeanCache (Gill et al., 2025) exemplifies a user-side semantic caching system tailored for LLM-powered web services. When a query is submitted to a MeanCache-enabled LLM service, the system computes an embedding vector for the query. This embedding is then compared against cached query embeddings using cosine similarity to identify potential matches. For each semantically similar query identified in the cache, MeanCache analyzes its associated context chain, comparing it with the dialogue history of the current query. If a match is found—that is, a cached query exhibits both semantic similarity and a congruent context chain—the corresponding response is retrieved from the local cache and served to the user. Otherwise, MeanCache forwards the query to the LLM service to elicit a response; this response, along with the query and its embedding, is subsequently stored in the cache. To preserve user privacy, MeanCache employs Federated Learning (FL) to train a compact embedding model, thereby obviating the need to store user data on a central server. The embedding model is trained using a multi-task learning approach that incorporates both contrastive loss and multiple-negatives ranking loss. This process also involves determining an optimal cosine similarity threshold, denoted as $\tau$, to strike a balance between true positives (cache hits) and false positives. To mitigate storage and computational overheads, MeanCache applies Principal Component Analysis (PCA) for dimensionality reduction of the embedding vectors. Empirical studies indicate that approximately 31% of user queries exhibit similarity to previously submitted ones, underscoring the potential of user-side caching to significantly reduce LLM inference costs.
@@ 6.3.3 Quantization
  • Weight-only Quantization. In weight-only quantization, models like ZeroQuant-V2 (Yao et al., 2023b) apply uniform quantization to OPT and BLOOM, finding that 8-bit weights with 16-bit activations preserve performance while 4-bit weights cause significant degradation. They introduce Low-Rank Compensation (LoRC) to approximate quantization error with a storage-efficient low-rank matrix. GLM-130B (Zeng et al., 2022) demonstrates successful 4-bit quantization using row-wise symmetric quantization, attributed to its well-shaped weight distributions. Non-uniform methods include LUT-GEMM/nuQmm (Park et al., 2022), which extends binary-coding quantization with bias terms and group-wise quantization. SqueezeLLM (Kim et al., 2023) uses sensitivity-based k-means centroids for non-uniform quantization, optimizing for memory-bound LLM inference. QLoRA (Dettmers et al., 2023) introduces NormalFormat datatype based on Quantile Quantization, normalizing weights to standard normal distribution. GPTQ (Frantar et al., 2023) builds on Optimal Brain Quantization, quantizing rows in parallel with lazy batch updates and Cholesky reformulation for stability. QuIP (Chee et al., 2023) defines optimal adaptive rounding methods using LDL decomposition and incoherence processing, achieving viable 2-bit quantization. AWQ (Lin et al., 2023) preserves significant activation channels in FP16 while quantizing others, using activation-aware scaling to better represent salient weights.
  • Weight+Activation Quantization. For weight+activation quantization, LLM.int8 (Dettmers et al., 2022) identifies systematic outliers in activations, preserving outlier dimensions in high-precision while quantizing average values to INT8. RPTQ (Yuan et al., 2023) clusters and reorders activation dimensions based on min/max values, fusing operations to reduce latency. Low-bit floating-point formats like FP4/FP8 are explored in MoFQ (Zhang et al., 2023b) and ZeroQuant-FP (Wu et al., 2023), providing better handling of activation outliers than integer formats. SmoothQuant (Xiao et al., 2023a) suppresses outliers by scaling down activation dimensions and scaling up corresponding weight dimensions with a controllable hyperparameter. Outlier Suppression (Wei et al., 2022b) uses inverse gamma values from LayerNorm as scaling factors, while Outlier Suppression+ (Wei et al., 2023) adds shifting factors to remove asymmetry in activations. FPTQ (Li et al., 2023a) introduces logarithmic activation equalization for non-linear moderation of activation distributions. OmniQuant (Shao et al., 2023a) learns optimal clipping ranges for extreme values, and QLLM (Liu et al., 2023) splits outlier channels into sub-channels while merging similar ones.

### v2

@@ 6.1 Background
AI-powered search applications have experienced a substantial increase in user traffic. For example, prominent platforms such as Perplexity are reported to manage over 10 million daily visits [perplexityai]. The core technology driving these applications consists of LLMs, which predominantly utilize the Transformer architecture [DBLP:conf/nips/VaswaniSPUJGKP17]. Although the Transformer model has demonstrated notable effectiveness, its computational complexity, which scales quadratically with the input sequence length, introduces significant challenges during inference. Moreover, in adherence to scaling laws [scalinglaw:arxiv20:Kaplan], LLMs have evolved to encompass an exceptionally large number of parameters. For instance, models such as LLaMA-4, DeepSeek-V3, and Qwen-3 possess parameter counts ranging from 100 billion to 1,000 billion, thereby requiring considerable storage capacity.
The computational requirements of LLMs necessitate considerable hardware resources, and their inference latency constitutes a significant portion of the overall response time in AI search applications. For example, a state-of-the-art LLM for AI search, DeepSeek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability], which reportedly incurs high deployment costs (requiring 32 GPUs, each with 80 GiB memory per instance) and suffers from latency bottlenecks (a first token latency of 2.91 seconds and a decoding speed of approximately 30 tokens/second). Search applications are inherently latency-sensitive, as users typically have limited tolerance for waiting times.
@@ 6.2.1 Local Attention
The internal computations of LLMs are predominantly characterized by matrix multiplications involving input context tokens and model parameter matrices. A prominent strategy to mitigate this computational burden is to restrict the scope of input tokens attended to by each token, a technique termed local attention. It is well-established that the computational complexity of the standard full attention mechanism exhibits quadratic scaling ($O(N^{2})$) with respect to the input sequence length $N$. However, a substantial body of research [xiao2023efficient] suggests that full attention is not always necessary, as only a subset of contextual tokens typically exerts a significant influence on the model’s output. Consequently, local attention mechanisms can often serve as an effective substitute for full attention.
@@
While various lightweight attention mechanisms have been proposed, such as linear attention [minimax2025minimax01scalingfoundationmodels], RWKV [peng2023rwkvreinventingrnnstransformer], and Mamba [gu2024mambalineartimesequencemodeling], their adoption typically requires training models de novo, incurring substantial costs. In contrast, local attention techniques can often be integrated into pre-trained full-attention models with minimal overhead. This adaptation can be accomplished either through lightweight fine-tuning [zhang2025lighttransfer] or, in some cases, applied directly at inference time without any retraining, as demonstrated by methods like StreamingLLM [xiao2023efficient].
@@
Recent works have explored various strategies to adapt pre-trained models. For instance, LightTransfer [zhang2025lighttransfer] addresses the high computational cost of full attention by identifying and modifying “lazy layers” where attention patterns are highly localized. Its innovation lies in selectively replacing full attention in these layers with a more efficient mechanism like streaming attention. This approach significantly reduces the KV cache footprint and computational load while preserving long-context capabilities, and can be applied either training-free or with minimal fine-tuning.
@@
More advanced approaches aim to linearize pre-trained Transformers to achieve linear-time inference. A key challenge in this conversion is integrating recurrent mechanisms, such as gates, without costly retraining. Liger [lan2025liger] tackles this problem by repurposing existing weights from the pre-trained model’s key projection to construct these gating mechanisms, thereby avoiding new parameters. Its hybrid attention, combining local sliding window attention with a linear recurrent model, effectively balances local context processing with long-range dependency modeling.
[…]
@@
Furthermore, models adapted with local attention can leverage existing, highly optimized attention libraries, such as FlashAttention [dao2022flashattentionfastmemoryefficientexact], during deployment, thereby ensuring computational efficiency in practice.
@@ 6.2.2 Model Pruning
Recent works have developed various innovative approaches to structured pruning. For instance, Layer Collapse (LaCo) [yang2024lacolargelanguagemodel] introduces a training-free method that directly removes entire layers from a pre-trained model. Its core innovation is the Reserving-Differences-while-Seeking-Common (RDSC) Layer Merge principle, which calculates the parameter differences between adjacent layers and merges this differential into a subsequent layer. This “collapses” layers, effectively pruning them while preserving the model’s overall function. A key advantage of LaCo is its ability to remove 30% to 50% of layers while maintaining performance and, crucially, preserving the model’s internal architecture. This allows the pruned model to serve as a drop-in replacement in existing systems without code modifications.
[…]
Building on classic compression frameworks, SlimGPT [ling2024slimgptlayerwisestructuredpruning] adapts the Optimal Brain Surgeon (OBS) technique for the unique challenges of structured pruning in LLMs. To overcome the limitations of OBS, SlimGPT introduces a Batched Greedy Pruning algorithm that enables low-cost and rapid structured pruning by processing groups of parameters. It employs a grouped Cholesky decomposition for attention heads and a dynamic group size for FFNs to achieve near-optimal pruning efficiently. A key advantage of SlimGPT is its accessibility: it is a task-agnostic method requiring only a small calibration dataset and can complete the entire pruning process in about an hour on a single GPU.
Another line of work targets semi-structured pruning, which is critical for hardware acceleration but can often degrade model accuracy. A novel plug-and-play post-training pruning (PTP) method [zhang2024plug] addresses this by introducing two key components. First, a new metric, Relative Importance and Activation (RIA), jointly considers weight magnitude and activation data to prevent the inadvertent removal of important weight channels, a drawback of previous metrics. Second, it employs an efficient channel permutation algorithm that reorganizes input channels to maximally preserve important weights before enforcing the N:M sparsity pattern. This training-free approach is scalable to models over 70B parameters and makes sparse LLM inference more practical by converting models to a hardware-friendly N:M format with minimal performance loss.
@@ 6.3.1 Output Length Reduction
  • Prompt-based Length Reduction. Some techniques leverage prompting to curtail output length. For example, TALE-EP [han2024token] estimates token budgets by first prompting the LLM to determine a reasonable token limit for a given task, then incorporating this estimate into a constraint prompt that guides the model to generate concise yet accurate responses. Chain-of-Draft (CoD) [xu2025chain] addresses the verbosity common in step-by-step reasoning by instructing models to “Think step by step, but only keep a minimum draft for each thinking step, with at most five words,” thereby preserving essential intermediate reasoning while significantly reducing token usage. Lee et al. [lee2025llmscompresschainofthoughttoken] conducted a systematic study of different compression instructions, revealing a universal trade-off between reasoning length and accuracy, and proposing the concept of “token complexity” as the minimum number of tokens required for successful problem-solving.
@@
  • Training-based Length Reduction. Other methods achieve output length reduction through modifications during the training phase. Demystifying [yeo2025demystifying] implements a Cosine Reward function based on Dirichlet principles, incorporating an “exceed length penalty” to control Chain-of-Thought (CoT) reasoning length while maintaining performance. Aggarwal et al. [aggarwal2025l1] modify the training data by incorporating explicit length constraint instructions, such as “Think for N tokens,” before applying policy optimization to pre-trained reasoning models. O1-Pruner [luo2025o1] introduces a Length-Harmonizing Reward that calculates the ratio between reference model and predicted CoT lengths, incorporating accuracy constraints to prevent performance degradation while shortening reasoning processes. DAST [shen2025dast] employs SimPO to fine-tune models using a specially constructed length-preference dataset based on a self-defined token-length budget measurement. Kimi [team2025kimi] incorporates a length penalty into its policy optimization approach to improve long CoT activations and facilitate model merging.
@@
  • Compression of Intermediate States. A distinct category of techniques focuses on compressing intermediate data generated during the model’s inference process. For instance, Coconut (Chain of Continuous Thought) [hao2024training] treats final-layer hidden states as “continuous thought” to replace discrete tokens, reusing these hidden states as subsequent input embeddings and gradually adding latent CoT tokens during training, thereby improving accuracy while reducing intermediate “thinking” tokens. CODI [shen2025codi] employs self-distillation where the model serves as both teacher and student, jointly learning explicit and implicit CoT while aligning hidden activations, enabling internal reasoning without generating explicit tokens. CCOT [cheng2024compressed] condenses long reasoning into short, contentful contemplation tokens by precomputing full CoT, selecting important hidden states, and training LoRA modules to predict and decode these key tokens.
@@ 6.3.2 Semantic Caching
AI search applications frequently encounter a high degree of semantic similarity among queries submitted by different users. Consequently, a considerable fraction of distinct input queries can yield identical or substantially similar answers (e.g., “Tell me about the history of Hawaii” versus “What is the history of Hawaii?”). This redundancy is leveraged by various optimization strategies, including query caching, result caching, and query rewriting [gill2025meancacheusercentricsemanticcaching].
For instance, MeanCache [gill2025meancacheusercentricsemanticcaching] exemplifies a user-side semantic caching system tailored for LLM-powered web services. When a query is submitted to a MeanCache-enabled LLM service, the system computes an embedding vector for the query. This embedding is then compared against cached query embeddings using cosine similarity to identify potential matches. For each semantically similar query identified in the cache, MeanCache analyzes its associated context chain, comparing it with the dialogue history of the current query. If a match is found—that is, a cached query exhibits both semantic similarity and a congruent context chain—the corresponding response is retrieved from the local cache and served to the user. Otherwise, MeanCache forwards the query to the LLM service to elicit a response; this response, along with the query and its embedding, is subsequently stored in the cache. To preserve user privacy, MeanCache employs Federated Learning (FL) to train a compact embedding model, thereby obviating the need to store user data on a central server. The embedding model is trained using a multi-task learning approach that incorporates both contrastive loss and multiple-negatives ranking loss. This process also involves determining an optimal cosine similarity threshold, denoted as $\tau$, to strike a balance between true positives (cache hits) and false positives. To mitigate storage and computational overheads, MeanCache applies Principal Component Analysis (PCA) for dimensionality reduction of the embedding vectors. Empirical studies indicate that approximately 31% of user queries exhibit similarity to previously submitted ones, underscoring the potential of user-side caching to significantly reduce LLM inference costs.
@@ 6.3.3 Quantization
  • Weight-only Quantization. In weight-only quantization, models like ZeroQuant-V2 [yao2023zeroquantv2] apply uniform quantization to OPT and BLOOM, finding that 8-bit weights with 16-bit activations preserve performance while 4-bit weights cause significant degradation. They introduce Low-Rank Compensation (LoRC) to approximate quantization error with a storage-efficient low-rank matrix. GLM-130B [zeng2022glm] demonstrates successful 4-bit quantization using row-wise symmetric quantization, attributed to its well-shaped weight distributions. Non-uniform methods include LUT-GEMM/nuQmm [park2022nuqmm], which extends binary-coding quantization with bias terms and group-wise quantization. SqueezeLLM [kim2023squeezellm] uses sensitivity-based k-means centroids for non-uniform quantization, optimizing for memory-bound LLM inference. QLoRA [dettmers2023qlora] introduces NormalFormat datatype based on Quantile Quantization, normalizing weights to standard normal distribution. GPTQ [frantar2023gptq] builds on Optimal Brain Quantization, quantizing rows in parallel with lazy batch updates and Cholesky reformulation for stability. QuIP [chee2023quip] defines optimal adaptive rounding methods using LDL decomposition and incoherence processing, achieving viable 2-bit quantization. AWQ [lin2023awq] preserves significant activation channels in FP16 while quantizing others, using activation-aware scaling to better represent salient weights.
  • Weight+Activation Quantization. For weight+activation quantization, LLM.int8 [dettmers2022llm] identifies systematic outliers in activations, preserving outlier dimensions in high-precision while quantizing average values to INT8. RPTQ [yuan2023rptq] clusters and reorders activation dimensions based on min/max values, fusing operations to reduce latency. Low-bit floating-point formats like FP4/FP8 are explored in MoFQ [zhang2023integer] and ZeroQuant-FP [wu2023zeroquant], providing better handling of activation outliers than integer formats. SmoothQuant [xiao2023smoothquant] suppresses outliers by scaling down activation dimensions and scaling up corresponding weight dimensions with a controllable hyperparameter. Outlier Suppression [wei2022outlier] uses inverse gamma values from LayerNorm as scaling factors, while Outlier Suppression+ [wei2023outlier] adds shifting factors to remove asymmetry in activations. FPTQ [li2023fptq] introduces logarithmic activation equalization for non-linear moderation of activation distributions. OmniQuant [shao2023omniquant] learns optimal clipping ranges for extreme values, and QLLM [liu2023qllm] splits outlier channels into sub-channels while merging similar ones.